OpenStudy (shortypants):

The personnel expenses were ​$8,700,000 in 2014. Over the previous 5​ years, the personnel expenses increased at a rate of 3.4​% annually. ​a) Assuming that this rate​ continues, write an equation describing personnel​ costs, C(t), in millions of​ dollars, where t=0 corresponds to 2014. ​b) Sketch a graph of this function up to the year 2029 ​(t=15​). ​c) What are your projected personnel costs in the year 2016​?

OpenStudy (shortypants):

The personnel expenses were ​$8,700,000 in 2014. Over the previous 5​ years, the personnel expenses increased at a rate of 3.4​% annually. ​a) Assuming that this rate​ continues, write an equation describing personnel​ costs, C(t), in millions of​ dollars, where t=0 corresponds to 2014. ​b) Sketch a graph of this function up to the year 2029 ​(t=15​). ​c) What are your projected personnel costs in the year 2016​?

OpenStudy (evoker):

Use the general exponential growth equation if you have it.

OpenStudy (shortypants):

alright i came out with C(t)=8700000(1+0.034)^16

OpenStudy (shortypants):

1.48541523106255409245115793655302259965918304468992 × 10^7 was my result on the calculator

OpenStudy (shortypants):

so 14854152.3106 million?

OpenStudy (evoker):

Hmm is the for b, as they said you want to use 15.

OpenStudy (evoker):

but otherwise you have the right equation

OpenStudy (shortypants):

yea 16 is for part c

OpenStudy (shortypants):

15 is a picture of a graph

OpenStudy (evoker):

actually 2 for c

OpenStudy (evoker):

your comparing to 2014, 2016 is 2 years later.

OpenStudy (shortypants):

oh

OpenStudy (shortypants):

plugged in 2 and got 9301657.2 You can search for any mathematical expression, using functions such as: sin, cos, sqrt, etc. You can find a complete list of functions here. RadDeg x! Inv sin ln π cos log e tan √ Ans EXP xy ( ) % AC 7 8 9 ÷ 4 5 6 × 1 2 3 − 0 . = +

OpenStudy (shortypants):

oops

OpenStudy (shortypants):

9301657.2

OpenStudy (evoker):

no problem, yeah I got the same result

OpenStudy (shortypants):

would C(t)=8700000(1+.034)^t be right?

OpenStudy (shortypants):

for A

OpenStudy (evoker):

Yes that is correct you could collapse the parenthesis to 1.034 if you wanted

OpenStudy (shortypants):

hmm i not sure what way the online program will say is correct

OpenStudy (shortypants):

thanks for your help!

OpenStudy (evoker):

no problem have a good night.